OpenStudy (anonymous):

If I am finding dy/dx it is the derivative of y/derivative of x, d^2y/dx^2 it is derivative of (dy/dx)/derivative of x, and d^3y/dx^3 would be derivative of (d^2y/dx^2)/derivative of x and so on and so on? keeping the denominator the derivative of x each time?

OpenStudy (anonymous):

dy/dx means the derivative of y with respect to x.

OpenStudy (anonymous):

And the correct Liebnitz notation is: First derivative: \[\frac{ dy }{ dx }\] Second derivative: \[\frac{ d^2y }{ dx^2 }\] Third derivative: \[\frac{ d^3y }{ dx^3 }\]

OpenStudy (anonymous):

but when working out the third derivative there on top would be the derifvative of the second derivative / by the derivative of the original x, yes?

OpenStudy (anonymous):

Third derivative of y is: \[\frac{ d }{ dx }(\frac{ d }{ dx }(\frac{ dy }{ dx }))\] Is this what you're asking?

OpenStudy (anonymous):

it just looks like each time you take the derivative of the one before and divide it by the derivative of the orginal x, like the d^3y = (d^2y/dx^2)/(dx/dt)

OpenStudy (anonymous):

Don't think of /dx as division! It's a notation, not a ratio.

OpenStudy (anonymous):

Want to try an easy example?

OpenStudy (anonymous):

yeah bud im down

OpenStudy (anonymous):

\[y = x^6\] \[\frac{ dy }{ dx } = 6x^5\] \[\frac{ d^2y }{ dx^2 } = \frac{ d }{ dx }(\frac{ dy }{ dx }) = 30x^4\]

OpenStudy (anonymous):

Are you following so far?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Cool. Try finding the third derivative.

OpenStudy (anonymous):

120x^3 right?

OpenStudy (anonymous):

yeap!

OpenStudy (anonymous):

but what if it is y= something and x=something else

OpenStudy (anonymous):

I'll make some more examples and you can try them :)

OpenStudy (anonymous):

like y=t^2) and x= t^4

OpenStudy (anonymous):

thanks for all the help my friend, got a cal 2 test in the morning and trying to sure things up

OpenStudy (anonymous):

try this: \[\frac{ d }{ dr }(r^3 + 2r^2)\]

OpenStudy (anonymous):

Wait so like: \[\frac{ d(t^2) }{ d(t^4) } ???\]

OpenStudy (anonymous):

3r^2 + 6r^2, but i am trying to find it when a given is y= and a x=

OpenStudy (anonymous):

differectial and parametric form it says

OpenStudy (anonymous):

I don't understand. You're trying to differentiate with respect to a function?

OpenStudy (anonymous):

i'm not sure of the technical way of saying it but asks for the parametric form of a third dericative, it gives a y= and a x= then for the second derivative you take the derivative of y/derivative of x then for the third i think it is the derviative of the second / the derivative of x=

OpenStudy (anonymous):

it not just the derivative of a y= but how to find a parametric formula for the second or third dericative when given a y= and a x=

OpenStudy (anonymous):

Is it perhaps: x = f(t), y = g(t)? And they want dy/dx? Ifso, you use the chain rule: (dy/dt)/(dx/dt)

OpenStudy (anonymous):

the questions read similar to this: For the parametric equations x=t +1 and y= t^2 +3t, find (d^2y)/(dx^2)

OpenStudy (anonymous):

Note that the higher order derivatives aren't so simple. \[\frac{ d^2y }{ dx^2 } = \frac{ \frac{ d }{ dt }(\frac{ dy }{ dx }) }{ \frac{ dx }{ dt } }\]

OpenStudy (anonymous):

looking at the notation and working backwards it looks like the numerator changes each tome to the next derivative while the denomenator stays the same no matter what degree you derive to?

OpenStudy (anonymous):

yeah thats it, to the third is it not the answer to that derived/ derivative of x=?

OpenStudy (anonymous):

I might not have to do to the third I have not checked, but I couldnt find anything in the book about it

OpenStudy (anonymous):

im on some mirc channels too, I might swing in there and ask too. I appreciate all the help with this and I am very thankfull for your time my friend. Thank you again, and have a good one.

OpenStudy (anonymous):

It's a pain to type but this site has a definition of the third derivative: http://17calculus.com/calc06-parametrics-calculus.php

OpenStudy (anonymous):

sweet thank you my friend

OpenStudy (anonymous):

Glad to help. Cheers!